Suppose the demand curve for a product is given by Q = 300 – 2P + 4I, where I is average income measured in thousands of dollars. The supply curve is Q = 3P – 50.

a. If I = 30, find the market clearing price and quantity for the product.

b. If I = 60, find the market clearing price and quantity for the product.

c. Draw a graph to illustrate your answers.



1
Expert's answer
2021-11-18T10:40:08-0500

a. Given that "I=30"

Then the demand curve "Q = 300 \u2013 2P + 4(30)"

or "Q = 420-2P"


To find marketing clearing price, demand should be equal to supply

"Q_d =Q_s"


"420\u22122P = 3P \u2013 50"


"5P = 470"


"P = 94"


Therefore


"Q=420\u22122(94)"


"Q=420- 188"


"Q =232"


b. Given that "I=60"

Then the demand curve "Q = 300 \u2013 2P + 4(60)"

or "Q = 540-2P"


To find marketing clearing price, demand should be equal to supply

"Q_d =Q_s"


"540\u22122P = 3P \u2013 50"


"5P = 590"


"P = 118"


Therefore


"Q=590\u22122(118)"


"Q=590- 236"


"Q =354"


c.